# On finite dimensional regular gradings

### v2

@@ 5. An explicit computation of the graded codimensions
In this section, we will use Theorem 38 in order to explicitly calculate the codimension sequence of a finite dimensional $G$-graded algebra whose regular decomposition is minimal and it has complete support. In particular, it will be shown that the graded PI-exponent of such algebras coincides with the ordinary PI-exponent.
@@
    exists and is a non-negative integer [2, Theorem 2.3].
@@
The proof of (i) is exactly the same as [5, Theorem 3.1], since the regularity of $B$ implies the existence of $a_{h_{1}}\in B_{h_{1}}$, …, $a_{h_{n}}\in B_{h_{n}}$ such that $a_{h_{1}}\cdots a_{h_{n}}\neq 0$. Similarly, the proof of (ii) is the same as the one given in [5, Corollary 3.2]. ∎
@@
Let $A$ be a finite dimensional $G$-graded algebra with regular grading and bicharacter $\beta$. Assume that the regular decomposition of $A$ is minimal and $A$ has complete support. Then $c_{n}^{G}(A)=|G|^{n}$, for all $n\in\mathbb{N}$. In particular, $\exp^{G}(A)=\exp(A)=|G|$.
@@ 6. General case
In this section, $A$ will denote a finite dimensional $G$-graded regular algebra with bicharacter $\beta$. As in the proof of Proposition 26 we write $A=D_{1}\oplus\cdots\oplus D_{k}+J(A)$, where $D_{1}\cong\cdots\cong D_{p}\cong K^{\alpha}G$, $\alpha\in H^{2}(G,K^{\ast})$, and $D_{i}\cong K^{\alpha_{i}}H_{i}$, where $H_{i}$ is a proper subgroup of $G$ and $\alpha_{i}\in H^{2}(H_{i},K^{\ast})$ for all $p<i\leq k$, $p<k$. We know by Remark 29, $D_{i}J(A)\neq 0$, for all $1\leq i\leq p$. Without loss of generality we can assume there exists $s\geq 0$, such that $D_{i}J(A)=0$, for all $p<i\leq p+s$. We will use the same notation as in the proof of Lemma 35.
We need the following technical lemmas.
Lemma 43.
If $A_{0}=E_{r_{1}}\oplus\cdots\oplus E_{r_{l}}+J(A_{0})$, then for each $1\leq i\leq k$, $(D_{i})_{0}=E_{r_{i}}$, for some $1\leq r_{i}\leq l$. In particular $d_{0}^{(i)}$ belongs to $E_{r_{i}}$, for all $1\leq i\leq k$.
Proof.
Since $H_{i}$ is finite, by [25, Theorem 2] $K^{\alpha_{i}}H_{i}$ is semisimple. Therefore the proof follows in the same way as the proof of item (1) in Lemma 35. ∎
Lemma 44.
If $J(A_{0})\neq 0$, then $s=0$.
Proof.
Suppose $s>0$ and take any $p<j\leq p+s$. Consider the $H_{j}$-graded algebra $V:=K^{\alpha_{j}}H_{j}+I$, where $I=\bigoplus_{h\in H_{j}}(J(A))_{h}$. Of course, by Remark 27, $V$ is a $H_{j}$-graded regular algebra (without unit), with bicharacter $\beta^{\prime}=\beta_{\mid_{H_{j}\times H_{j}}}$. By Lemma 43 we know that $(K^{\alpha_{j}}H_{j})_{0}=E_{r_{j}}$. Since $E_{j}$ is a copy of $K$, II can be seen as a vector space over $E_{j}$, in particular the algebra $V$ is a vector space over $E_{j}$. Considering $E_{j}$ with the trivial $G$-grading, for any $n\in\mathbb{N}$, by Theorem 41 we get
$c_{n}^{G}(V)=c_{n}^{G}(V\otimes_{E_{j}}E_{j})=|H_{j}|^{n}c_{n}^{G}(E_{j})=|H_{j}|^{n},$
therefore $\exp^{{H_{j}}}(V)=|H_{j}|$. Since $\exp^{{H_{j}}}(V)\leq|H_{j}|\exp(V)$ we conclude $\exp(V)\geq 1$. Consequently, by construction, if $J(V)\neq 0$, then $(K^{\alpha_{j}}H_{j})I\neq 0$, that is a contradiction because $(K^{\alpha_{j}}H_{j})J(A)=0$ and $I\subseteq J(A)$. We conclude $J(V)=0$, i.e, $I=0$. In particular, $J(A)_{0}=0$, that is a contradiction because $J(A_{0})\neq 0$ and $J(A_{0})\subseteq J(A)_{0}$. ∎
Lemma 45.
If $J(A_{0})\neq 0$, then $E_{r_{i}}J(A_{0})\neq 0$, for all $1\leq i\leq k$.
Proof.
The proof follows from the same argument used in the proof of item (3) in Lemma 35, because as we have just seen in Lemma 44, $D_{i}J(A)\neq 0$, for all $1\leq i\leq k$. ∎
Having Lemma 43 at hand, the proof of item (2) of Lemma 35 is entirely independent of the assumptions of complete support and minimal regular decomposition. Therefore, the same arguments yield the following.
@@
Based on the results above, we have the following situation. The decomposition $A=D_{1}\oplus\cdots\oplus D_{k}+J(A)$ is such that:
@@
Let $A$ be a finite dimensional $G$-graded algebra with bicharacter $\beta$.
@@
    Suppose the regular decomposition of $A$ is minimal and $J(A_{0})\neq 0$. Then, there exist $\mathscr{V}$ a commutative algebra, and $\mathscr{W}$ a $G$-graded algebra which is not regular such that
@@
Thus, if we take $\mathscr{W}:=W_{p+1}\oplus\cdots\oplus W_{p+s}$, then the result follows. Now, by the definition of $\mathscr{W}$ and by the same argument used at the end of the proof of Lemma 36, we obtain
$J(\mathscr{W})_{0}=1_{E_{r_{p+1}}}J(A_{0})+\cdots+1_{E_{r_{k}}}J(A_{0})$
@@
$J(\mathscr{V})+J(\mathscr{W})_{0}=1_{E_{r_{1}}}J(A_{0})+\cdots+1_{E_{r_{k}}}J(A_{0})=J(A_{0})$
and this proves (i).
Let us prove (ii). Suppose $J(A_{0})=0$. We know that there exists $d_{g}^{(1)}\in D_{1}$, with some $g\in G$, such that $d_{g}^{(1)}z\neq 0$, for some homogeneous element $z\in(J(A))_{\kappa}$, $\kappa\in G$. In particular, by multiplication by $(d_{g}^{(1)})^{-1}$, we conclude $1_{E_{1}}z\neq 0$. If for some $h\in G$ we have $d_{h}^{(1)}z=0$, then $1_{E_{1}}z=0$, which is not the case. Therefore, $d_{h}^{(1)}z\neq 0$, for all $h\in G$, that implies $d_{-\kappa}^{(1)}z\neq 0$ whereas $d_{-\kappa}^{(1)}z\in J(A_{0})$ which is a contradiction. We conclude $J(A)=0$, so $A$ is a direct sum of twisted group algebras with $D_{1}\cong K^{\alpha}G$, where $\alpha\in H^{2}(G,K^{\ast})$. ∎
The following corollary is then immediate.
Corollary 49.
Let $A$ be a finite dimensional $G$-graded algebra whose regular decomposition is minimal. Then, there exists a finite dimensional $G$-graded regular algebra whose regular decomposition is minimal, has complete support and it is a direct summand of $A$.
@@
Let $A$ be a finite dimensional $G$-graded regular algebra with bicharacter $\beta$. Then, the regular decomposition of $A$ is minimal if and only if $\exp(A)=|G|$.
@@
By Theorem 10 we just need to show that $\exp(A)=|G|$ implies the regular decomposition of $A$ is minimal. Let us write $A=D_{1}\oplus\cdots\oplus D_{k}+J(A)$, where by Theorem 48, without loss of generality we can assume $D_{1}\cong K^{\alpha}G$, $\alpha\in H^{2}(G,K^{\ast})$, and $D_{i}\cong K^{\alpha_{i}}H_{i}$, where $H_{i}$ is a subgroup of $G$ and $\alpha_{i}\in H^{2}(H_{i},K^{\ast})$, for all $2\leq i\leq k$. Suppose $\exp(A)=|G|$ and the regular decomposition of $A$ is non-minimal. In particular, $\det M^{A}=0$. Given $i\neq j$, suppose $D_{i}J(A)D_{j}\neq 0$. Then, there exist homogeneous elements $d\in D_{i}$, $u\in J(A)$ and $d^{\prime}\in D_{j}$ such that $dud^{\prime}\neq 0$. On the other hand, since $A$ is $\beta$-commutative, there exists $\lambda\neq 0$ such that $dud^{\prime}=\lambda udd^{\prime}$ which implies $dud^{\prime}=0$ because $D_{i}D_{j}=0$. Therefore for all $1\leq i,j\leq k$ we have $D_{i}J(A)D_{j}=0$. By Yamazaki’s Theorem [19, Corollary 8.2.10], for all $1\leq i\leq k$, there exist $n_{i}$, $m_{i}\in\mathbb{N}$ such that
$D_{i}\cong\underbrace{M_{n_{i}}(K)\oplus\cdots\oplus M_{n_{i}}(K)}_{m_{i}}.$
In this way, if $J(A_{0})\neq 0$, then by Lemma 44, $D_{i}J(A)\neq 0$, for all $1\leq i\leq k$, and by the construction of the PI-exponent we get
(7) $|G|=\exp(A)=\max\{\dim M_{n_{i}}(K)\mid 1\leq i\leq k\}.$
It is worth noting that if $J(A_{0})=0$, then by Theorem 48 we have $J(A)=0$, i.e., $A$ is semisimple. Hence, by [18, Theorem 6.6.1], $\exp(A)=\max\{\dim M_{n_{i}}(K)\mid 1\leq i\leq k\}$. Thus, (7) holds for any finite dimensional regular algebra. Now take $1\leq t\leq k$ such that $\exp(A)=\dim M_{n_{t}}(K)=n_{t}^{2}$. Since $\exp(A)=|G|$, it follows that $n_{t}^{2}=|G|$ allowing us to conclude that $m_{t}=1$ and $D_{t}\cong K^{\alpha_{t}}G\cong M_{\sqrt{|G|}}(K)$. Because $M_{\sqrt{|G|}}(K)$ is central simple, it follows that $K^{\alpha_{t}}G$ is central simple too. Moreover, since $\alpha_{t}$ induces $\beta$, by Lemma 20, the regular decomposition of $D_{t}$ is minimal and jointly with Theorem 10 we get $\det M^{A}=\det M^{D_{t}}\neq 0$ which is a contradiction. ∎

### v3

@@ 5. An explicit computation of the graded codimensions
In this section, we will use Theorem 38 in order to explicitly calculate the codimension sequence of a finite dimensional $G$-graded algebra $A$ with bicharacter $\beta$, complete support, and minimal regular decomposition. Moreover, if we write $A=D_{1}\oplus\cdots\oplus D_{k}+J(A)$ , where $D_{i}$ is a $G$-graded simple subalgebra of $A$, then $D_{i}J(A)\neq 0$, for all $1\leq i\leq k$. In particular, it will be shown that the graded PI-exponent of such algebras coincides with the ordinary PI-exponent.
@@
    exists and is a non-negative integer [3, Theorem 2.3].
@@
The proof of (i) is exactly the same as [6, Theorem 3.1], since the regularity of $B$ implies the existence of $a_{h_{1}}\in B_{h_{1}}$, …, $a_{h_{n}}\in B_{h_{n}}$ such that $a_{h_{1}}\cdots a_{h_{n}}\neq 0$. Similarly, the proof of (ii) is the same as the one given in [6, Corollary 3.2]. ∎
@@
For any $n\in\mathbb{N}$, $c_{n}^{G}(A)=|G|^{n}$. In particular, $\exp^{G}(A)=\exp(A)=|G|$.
@@ 6. General case
In this section, we consider the general case, namely, when the algebra $A$ does not necessarily have complete support. As in the proof of Proposition 26 we write $A=D_{1}\oplus\cdots\oplus D_{k}+J(A)$, where $D_{1}\cong\cdots\cong D_{p}\cong K^{\alpha}G$, $\alpha\in H^{2}(G,K^{\ast})$, and $D_{i}\cong K^{\alpha_{i}}H_{i}$, where $H_{i}$ is a proper subgroup of $G$ and $\alpha_{i}\in H^{2}(H_{i},K^{\ast})$ for all $p<i\leq k$. Also, we will assume $D_{i}J(A)\neq 0$ for all $1\leq i\leq k$.
Notice if $p=0$ then $A$ is just a $\beta$-commutative algebra which is not regular and if $p=k$ then $A$ has complete support. Given that our analysis will implicitly cover the first case, we can assume without loss of generality $1\leq p<k$.
From now on, we will use the same notation as in the proof of Lemma 35.
Notice the proof of item (2) of Lemma 35 is entirely independent of the assumptions of complete support. Therefore, the same arguments yield the following.
@@
Lemma 44.
If $A_{0}=E_{r_{1}}\oplus\cdots\oplus E_{r_{k}}+J(A_{0})$, then for each $1\leq i\leq k$, $(D_{i})_{0}=E_{r_{i}}$, for some $1\leq r_{i}\leq l$. In particular $d_{0}^{(i)}$ belongs to $E_{r_{i}}$, for all $1\leq i\leq k$.
Proof.
Since $(H_{i})_{0}\cong K$, the result follows as in item (1) of Lemma 35. ∎
Lemma 45.
If $J(A_{0})\neq 0$, then $E_{r_{i}}J(A_{0})\neq 0$, for all $1\leq i\leq p$.
Proof.
The proof follows from the same argument used in the proof of item (3) in Lemma 35. ∎
@@
Based on the results above, the decomposition $A=D_{1}\oplus\cdots\oplus D_{k}+J(A)$ satisfies the following:
  1. (1)
    $D_{i}J(A)\neq 0$, for all $1\leq i\leq k$.
@@
Remark 47.
Given $p+1\leq i\leq k$, we have $1_{E_{r_{i}}}J(A)\neq 0$. Indeed, since $D_{i}J(A)\neq 0$, there exists $d_{h}^{(i)}\in\mathcal{D}_{i}$ such that $d_{h}^{(i)}z\neq 0$, for some $z\in J(A)$. If $1_{E_{r_{i}}}z=0$, we would have
$d_{h}^{(i)}z=\alpha(h,0)^{-1}d_{h}^{(i)}1_{E_{r_{i}}}z=0$
that is a contradiction.
@@
    Suppose $J(A_{0})\neq 0$. Then, there exist $\mathscr{V}$ a commutative algebra, and $\mathscr{W}$ a $G$-graded algebra which is not regular such that
@@
Thus, if we take $\mathscr{W}:=W_{p+1}\oplus\cdots\oplus W_{k}$, then the result follows. Now, by the definition of $\mathscr{W}$ and by the same argument used at the end of the proof of Lemma 36, we obtain
$J(\mathscr{W}_{0})=1_{E_{r_{p+1}}}J(A_{0})+\cdots+1_{E_{r_{k}}}J(A_{0})$
@@
$J(\mathscr{V})+J(\mathscr{W}_{0})=1_{E_{r_{1}}}J(A_{0})+\cdots+1_{E_{r_{k}}}J(A_{0})=J(A_{0})$
and this proves (i). We observe at this point that, for some $p+1\leq i\leq k$, it may occur that $1_{E_{r_{i}}}J(A_{0})=0$.
Let us prove (ii). Suppose $J(A_{0})=0$. We know that there exists $d_{g}^{(1)}\in D_{1}$, with some $g\in G$, such that $d_{g}^{(1)}z\neq 0$, for some homogeneous element $z\in(J(A))_{\kappa}$, $\kappa\in G$. In particular, by multiplication by $(d_{g}^{(1)})^{-1}$, we conclude $1_{E_{r_{1}}}z\neq 0$. If for some $h\in G$ we have $d_{h}^{(1)}z=0$, then $1_{E_{1}}z=0$, which is not the case. Therefore, $d_{h}^{(1)}z\neq 0$, for all $h\in G$, that implies $0\neq d_{-\kappa}^{(1)}z\in A_{0}\cap J(A_{0})=J(A_{0})$, which is a contradiction. We conclude $J(A)=0$, so $A$ is a direct sum of twisted group algebras with $D_{1}\cong K^{\alpha}G$, where $\alpha\in H^{2}(G,K^{\ast})$. ∎
Remark 49.
The hypothesis that $D_{i}J(A)\neq 0$ cannot be removed in the Theorem 48. Indeed, considering the Example 28, in that regular grading we have $J(\mathcal{A}_{0})=0$, but $J(\mathcal{A})\neq 0$.
With the results of this section, together with the discussion preceding Example 28, we obtain the main result of this work:
Theorem 50.
Let $\mathcal{B}$ be a finite-dimensional $G$-graded regular algebra with bicharacter $\beta$. Then there exist $\mathcal{A}$ and $\mathcal{S}$, both $G$-graded subalgebras of $\mathfrak{B}$, with $\mathcal{S}=K^{\alpha_{1}}Q_{1}\oplus\cdots\oplus K^{\alpha_{t}}Q_{t}$, $\mathcal{B}=\mathcal{A}\oplus\mathcal{S}$ (direct sum as algebras) and $J(\mathcal{B})=J(\mathcal{A})$, such that the following hold:
  1. (1)
    If $J(\mathcal{B})\neq 0$, then $\mathcal{A}\neq 0$ and the product of any graded simple component of $\mathcal{A}$ with the Jacobson radical of $\mathcal{B}$ is nonzero.
  2. (2)
    If $\mathcal{A}$ has non-graded simple components isomorphic to some $K^{\alpha}G$, then $\mathcal{S}$ must contain at least one isomorphic to $K^{\alpha}G$, and $\mathcal{A}$ is a $\beta$-commutative algebra that is not a regular $G$-grading.
  3. (3)
    If $\mathcal{A}$ has a graded simple component isomorphic to $K^{\alpha}G$, then we have the options:
    1. (i)
      If $J(\mathcal{B}_{0})\neq 0$, there exist $\mathscr{V}$, a commutative algebra, and $\mathscr{W}$, a $G$-graded algebra which is not regular, such that
      $\mathcal{A}\cong(K^{\alpha}G\otimes\mathscr{V})+\mathscr{W}\quad(\text{as $G$-graded algebras})$
      and $J(\mathscr{V})+J(\mathscr{W}_{0})=J(\mathcal{B}_{0})$.
    2. (ii)
      If $J(\mathcal{B}_{0})=0$, then $J(\mathcal{B})=0$. In particular, $\mathcal{A}$ is semisimple.
@@
Let $\mathcal{B}$ be a finite dimensional $G$-graded regular algebra with bicharacter $\beta$. Then, the regular decomposition of $\mathcal{B}$ is minimal if and only if $\exp(\mathcal{B})=|G|$.
@@
By Theorem 10 we just need to show that $\exp(\mathcal{B})=|G|$ implies the regular decomposition of $\mathcal{B}$ is minimal. Since $T(\mathcal{B})=T(K^{\alpha}G)$, for some $\alpha\in H^{2}(G,K^{\ast})$, it is enough to show that $\exp(K^{\alpha}G)=|G|$ implies that $K^{\alpha}G$ is a central simple algebra. By Yamazaki’s Theorem [20, Corollary 8.2.10], there exist $n$, $m\in\mathbb{N}$ such that
$K^{\alpha}G\cong\underbrace{M_{n}(K)\oplus\cdots\oplus M_{n}(K)}_{m}.$
and by the construction of the PI-exponent we have $|G|=\exp(A)=\dim M_{n}(K)$. Since $\dim K^{\alpha}G=|G|$, we conclude $K^{\alpha}G\cong M_{n}(K)$ are we are done. ∎